Session: Nov 1998 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Thu Nov 26 13:56:29 MET 1998) Minor failures
No useful recording from BBC07 throughout the experiment.

Effelsberg (dgraham at Thu Nov 26 14:05:30 MET 1998) Success
vc01-06 active . RLMi looks strong, 12Jy peak of line at 1665. Not enough slewing time between source and 3c286.

Onsala (A. Polatidis at Thu Nov 26 17:39:45 MET 1998) Minor failures
No problems with the observations. Only VC01-VC04 are equiped with 0.5 MHz filters. There is no data for VC05-VC08, sorry about this.

Medicina (mariotti at Fri Nov 27 09:35:46 MET 1998) Minor failures
No local problem. The standard value of the local oscillator given prom the PI is not the proper one for the three lines be observed at the same time. The operator changed the LO value (the PI is aware of that). Medicina has 0.5 filters in the first six Vcs only, so the VC7,8 have no signal.

Jodrell (Alastair Gunn at Fri Nov 27 12:03:26 MET 1998) Minor failures
Telescope parked at 12:11 UT because of receiver power failure. Last twenty minutes lost. Filter from previous experiment was not removed so no signal likely to be found at lower edge line.

Noto (C.Trigilio at Mon Nov 30 11:13:04 MET 1998) Success

Westerbork (Tony Foley at Tue Dec 1 18:17:35 MET 1998) Minor failures
Only first 7 VCs had 0.5MHz filters. Polarizations were swapped. Scans after 10:59UT are out of Hour Angle range.

Calibration (Jean\-Francois Desmurs at Fri Dec 11 16:32:16 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: ee002a.antab. Read the comments in the file to get further information.